OpenStudy (anonymous):

evaluate the integral: 2x+5/x+1 dx

terenzreignz (terenzreignz):

Let's see... \[\huge \int \frac{2x+5}{x+1}dx\]

terenzreignz (terenzreignz):

What approaches have you tried?

OpenStudy (anonymous):

this will help: split it up into two integrals, 2x/x+1 and 5/x+1. use u substitution and manipulate u.

terenzreignz (terenzreignz):

No, @lilminiwin you do split the two integrals, but there is a much better way, methinks... I'll wait for @Juan07 's approach.

terenzreignz (terenzreignz):

Okay.... I'm alone~ No matter.... You split it like this... \[\Large \int \frac{2x+5}{x+1}dx = \int \frac{2x+2+3}{x+1}dx = \int \left[\frac{2x+2}{x+1}+\frac3{x+1}\right]dx\] and that's much easier.

OpenStudy (anonymous):

I'll concede that that was slightly easier.

terenzreignz (terenzreignz):

Still confused, @Juan07 ? How about I do one more small step... \[\Large = \int \left[\frac{2(x+1)}{x+1}+\frac3{x+1}\right]dx\]

OpenStudy (anonymous):

ok thanks for your help

OpenStudy (anonymous):

:D

terenzreignz (terenzreignz):

No problem.